a store manager wishes to reduce the price on her fresh ground coffee by mixing two grades. if she has 50 pounds of coffee which sells for $10 per pound, how much coffee worth $6 per pound must she mix with it so that she can sell the final mixture for $8.50 per pound?
Answer by mananth(16946)

a store manager wishes to reduce the price on her fresh ground coffee by mixing two grades. if she has 50 pounds of coffee which sells for $10 per pound, how much coffee worth $6 per pound must she mix with it so that she can sell the final mixture for $8.50 per pound?
...
50 pounds -------- $10 per pound,
x pounds --------- $6 per pound
x+50 pounds------- $8.50 per pound
..
mixture equation
50*10+6*x=8.50(x+50)
500+6x=8.50x+425
8.50x-6x=500-425
2.50x=75
x= 30 lb.of $6 coffee.
...
m.ananth@hotmail.ca